OpenStudy (anonymous):

Best answer will be given Write the quadratic variation equation for the relationship: y varies directly with x^2 and y = 36 when x = 3.

OpenStudy (anonymous):

Since y varies directly with x^2, the equation is y = kx^2 (k is the constant) Since y=36 when x=3, substituting these values gives 36 = k(9) k=4 So the quadratic variation equation is y=4x^2